OpenStudy (anonymous):

what is the equation of a circle whose center is (-5,3) and passes through the point (-1,3)?

jigglypuff314 (jigglypuff314):

circle formula: (x - h)^2 + (y - k)^2 = r^2 when center is (h, k) and radius (r) is the distance between (-5, 3) and (-1, 3)

OpenStudy (anonymous):

this is a really dumb question but do we have a radius yet?

jigglypuff314 (jigglypuff314):

what do you mean "do we have a radius yet?" you can get the radius by finding the distance between (-5, 3) and (-1, 3)

OpenStudy (anonymous):

oh okay sorry

jigglypuff314 (jigglypuff314):

it's alright :) then once you have all your values, you can plug it into the circle formula ^.^

OpenStudy (anonymous):

thank you:)) I just honestly have no idea what im doing here!!

jigglypuff314 (jigglypuff314):

then once you think you've got your formula could you post it so I could help check it over for you? :)

OpenStudy (anonymous):

wait- so I got that the radius was the square root of 16,but now what numbers do I plug into the circle formula?

jigglypuff314 (jigglypuff314):

circle formula: (x - h)^2 + (y - k)^2 = r^2 when center is (h, k) so the center you have is (-5, 3) so (x-(-5))^2+(y-(3))^2=(√16)^2 which can be simplified

OpenStudy (anonymous):

ohhh!okay so I got (x+5)^2+(y-3)^2=16

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thank you!! would you mind if you checked out another one?(:

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

what are the coordinates of the center of a circle if the endpoints of its diameter are A(8,-4) and B(-3,2)?

jigglypuff314 (jigglypuff314):

find the midpoint of (8, -4) and (-3, 2)

OpenStudy (anonymous):

I got M=(2.5,-1)

jigglypuff314 (jigglypuff314):

yup :) so that's the center of the circle :D

OpenStudy (anonymous):

finally! geometry is so confusing

OpenStudy (anonymous):

:p

jigglypuff314 (jigglypuff314):

lol your welcome :)

OpenStudy (anonymous):

can I send you another one because I just realized I don't know how to do this whole worksheet???!

jigglypuff314 (jigglypuff314):

lol sure :)

OpenStudy (anonymous):

which set of equations represents two circles that have the same center? 1)x^2+(y+4)^2=16 and (+4)^2+y^2=16 2)x-7^2+y-2^2=16 and x+7^2+y+^2=25 3)x-2^2+y-5^2=16 and x-2^2=y-5^2=25

jigglypuff314 (jigglypuff314):

well knowing that circle formula: (x - h)^2 + (y - k)^2 = r^2 when center is (h, k) what values match? ^ ^

OpenStudy (anonymous):

1 ?

jigglypuff314 (jigglypuff314):

no because x^2+(y+4)^2=16 and (x+4)^2+y^2=16 ^ and ^ don't match ^ and ^

OpenStudy (anonymous):

2? but they don't have the same radius

OpenStudy (anonymous):

wait 3 actually

jigglypuff314 (jigglypuff314):

3 is correct :) because it was just asking for same center, not same radius

OpenStudy (anonymous):

ohhhh okay thank you! only one more!

OpenStudy (anonymous):

a circle has the equation (x-3)^2+(y+4)^2=10. find the coordinates of the center of the circle and the length of its radius

jigglypuff314 (jigglypuff314):

well knowing that circle formula: (x - h)^2 + (y - k)^2 = r^2 when center is (h, k) center would be ^ ^ and radius would be ^

OpenStudy (anonymous):

center would be (-3,4) and radius would be the square root of 10?

jigglypuff314 (jigglypuff314):

ah, that's where most people get it wrong see, in the equation (x - h)^2 + (y - k)^2 = r^2 when center is (h, k) see how that is a negative h? and a negative k? so it's more like (x + (-h))^2 + (y + (-k))^2 = r^2 so in (x-3)^2+(y+4)^2=10 think of it as (x+(-3))^2 +(y+(4))^2=10 so your center would be (3, -4)

OpenStudy (anonymous):

ughhh I was thinking that too but I wasn't sure- so you basically just have to do it backwards- like find out the original problem?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

thank you so much!! math is not my strong point as you can probably see lol

jigglypuff314 (jigglypuff314):

it's fine :) I glad I can help people out :P i gtg good night ^.^

OpenStudy (anonymous):

:)) goodnight!